a standard baseball diamond is a square 90 ft. on a side. find the distance of a throw made from the catcher approximately 3 ft. behind home plate in an attempt to throw out a runner trying to steal second base: a standard baseball diamond is a square 90 ft. on a side. find the distance of a throw made from the catcher approximately 3 ft. behind home plate in an attempt to throw out a runner trying to steal second base
Answer by nycsharkman(136)

A standard baseball diamond is a square 90 feet on a side. Find the distance of a throw made from the catcher approximately 3 feet behind home plate in an attempt to throw out a runner trying to steal second base.
Pythagorean theorem:
a^2 + b^2 = c^2
We are looking for c + 3 feet (the distance where the catcher is situated behind home plate).
We must first find c.
(90)^2 + (90)^2 = c^2
16200 = c^2
Take square root of both sides.
127.2792206 = c
We can round 127.2792206 to the nearest ones place.
Doing this we get:
127 feet = c
However, the catcher is 3 feet BEHIND home plate.
We add 3 feet to 127 feet and we get 130 feet.
Did you follow?